Enrollment in a particular class for the last six semesters had been 20, 30, 50, 40, 80 and 60. Develop a forecast of enrollment next semester using exponential smoothing with an alpha = 0.2 and use the naive method to start the exponential smoothing process (i.e., forecast for semester 2 = 20).
Answer #1

Solution:

In exponential smoothing,

F(t+1) = \alpha A(t) + (1-\alpha)F(t)

where,

F(t+1) = Forecast for the next period

A(t) = Actual demand for the current period

F(t) = Forecast for the current period

\alpha = Smoothing constant

\alpha = 0.2

(1-\alpha) = 0.8

The forecast of enrollment next semester using exponential smoothing is calculated in the table below:

From the above calculations, Forecast for 7th semester = 44.05